1. y = c1ex + c2 xex ; y” -2y’ + y = 0


2. 2. x2 + y2 = r2 ; y’ = - 𝑥 /y






1
Expert's answer
2021-09-09T08:05:09-0400

1. Let us solve the equation "y'' -2y' + y = 0." The characteristic equation "k^2-2k+1=0" is equivalent to "(k-1)^2=0," and hence it has the roots "k_1=k_2=1." Therefore, its general solution is indeed "y = C_1e^x + C_2 xe^x."


2. Let us solve the equation "y' = - \frac{𝑥}{y}" is equivalent to "\frac{dy}{dx} = - \frac{𝑥}{y}," and hence to "ydy=-xdx." It follows that "\int ydy=-\int xdx," and consequently "\frac{y^2}{2}=-\frac{x^2}{2}+C." Since "C=\frac{y^2}{2}+\frac{x^2}{2}\ge0," we conclude that "2C\ge 0," and hence "2C=r^2" for some "r." It follows that "\frac{r^2}{2}=\frac{y^2}{2}+\frac{x^2}{2}," and therefore the general solution of the differential equation is indeed of the form "x^2 + y^2 = r^2."